It has been suggested that rotating cylinders about 14.0 mi long and 3.94 mi in diameter be placed in space and used as colonies. What angular speed must such a cylinder have so that the centripetal acceleration at its surface equals the free-fall acceleration on Earth?


___rad/s


1
Expert's answer
2021-10-05T15:44:25-0400

Solution.

"a=9.8 m\/s^2;"

"d=3.94 mi=6340.815m;"

"a=\dfrac{v^2}{r}\implies v^2=ar;"

"v^2=9.8m/s^2\sdot3170.4=31070(m/s)^2;"

"v=\sqrt{31070}=176.27m/s;"

"w=\dfrac{v}{r}=0.056 rad/s;"

Answer: 0.056 rad/s.